OpenStudy (anonymous):

–8.7 = k + 6 A. k = –8.7 – 6 B. k = 6 – 8.7 C. k = 8.7 – (–6)

OpenStudy (anonymous):

Please help

sammixboo (sammixboo):

To solve this question you must subtract 6 on both sides of the equation -8.7 = k + 6 k + 6 = -8.7 k + 6 - 6 = -8.7 - 6 k = -8.7 - 6 So what is -8.7 - 6